### v1

# FEATURE ARTICLE Visualization for Trust in Machine Learning Revisited: The State of the Field in 2023

Abstract
Visualization for explainable and trustworthy machine learning remains one of the most important and heavily researched fields within information visualization and visual analytics with various application domains, such as medicine, finance, and bioinformatics. After our 2020 state-of-the-art report comprising 200 techniques, we have persistently collected peer-reviewed articles describing visualization techniques, categorized them based on the previously established categorization schema consisting of 119 categories, and provided the resulting collection of 542 techniques in an online survey browser. In this survey article, we present the updated findings of new analyses of this dataset as of fall 2023 and discuss trends, insights, and eight open challenges for using visualizations in machine learning. Our results corroborate the rapidly growing trend of visualization techniques for increasing trust in machine learning models in the past three years, with visualization found to help improve popular model explainability methods and check new deep learning architectures, for instance.
Trust in machine learning (ML) models is a major concern in leveraging these technologies for real-world applications. Yet, ML models are being deployed in different application fields, and their role in decision-making processes is growing rapidly. Fields such as healthcare and criminal justice increasingly depend on ML models to make irreversible decisions that impact human lives. However, the black-box nature of some ML models poses a threat to their adoption. Domain experts often hesitate to rely on ML models for high-risk decision-making, as the inability to understand their inner workings fosters mistrust.
In response to the outlined challenges, researchers in academia and industry have designed several innovative solutions. For example, Google’s Explainable Artificial Intelligence (AI) Cloud and Descriptive mAchine Learning EXplanations (DALEX) package aims to improve the collaboration among domain experts to address the challenges posed by the complexity of AI. Except for all the visualization techniques analyzed in this survey article, recent frameworks set a foundation for developing techniques that facilitate users in communicating and externalizing their trust explicitly across varied ML stages and in understanding the complexities of human and AI interactions. Other works bridge the significant gap between ML outputs and human cognition by promoting a cross-disciplinary approach and building a robust model designed to examine the sender’s explanation intention and its ensuing impact on the receiver’s perception.
We base this work upon the findings of our previously published survey articles and others that have stressed the need for visual analytics (VA) to improve trust and transparency in areas such as dimensionality reduction (DR), deep learning (DL), and ML in general. After our 2020 state-of-the-art report (STAR) comprising 200 techniques, we have been collecting peer-reviewed articles describing visualization techniques for enhancing trust in ML, categorizing them based on the previously established categorization schema of 18 groups and 119 categories in total, and providing the hand-curated compilation of 542 visualization techniques in an online survey browser available at:
Although this survey article follows the same styling and overall structure as our STAR published in 2020, here, we made a new analysis of the updated dataset and discuss the most recent findings that provide us with an overview of the trust in ML visualization field as of fall 2023, including insights about the visualization community. The contributions of this article are:
the 542 techniques from the past 15 years categorized in trust levels (TLs) of interactive ML (compared to 200 in the 2020 STAR);
the trends and category correlations found using topic, temporal, correlation, and pattern mining (new compared to the 2020 STAR) analyses;
the interactive survey browser that aids domain experts, ML experts, visualization researchers, and others in exploring the field’s literature; and
the eight open challenges in VA for increasing the trustworthiness of the ML process (new compared to the six challenges of the 2020 STAR).
Note:
() The data collection was completed in September 2023.
() This technique was found in one article’s Related Work section.
Note: Rows are conferences except if journals (‘J’) or workshops (‘W’).
In this survey article, we rely on the same literature search methodology and run an identical analysis as in our 2020 STAR. We also adopt the same general definition of trust by Lee and See: “the attitude that an agent will help achieve an individual’s goals in a situation characterized by uncertainty and vulnerability” towards a more detailed, multi-level model of trust in ML that consists of five trust levels: the raw data (TL1: source reliability and transparent collection process), the processed data (TL2: uncertainty awareness, equality/data bias, comparison of structures, guidance/recommendation, and outlier detection), the algorithm/learning method (TL3: familiarity, understanding/ explanation, debugging/diagnosis, refinement/steering, comparison, knowledgeability, and fairness), the concrete model(s) for a particular task (TL4: experience, in situ comparison, performance, what-if hypotheses, model bias, and model variance), and the evaluation and the subjective users’ expectations (TL5: agreement of colleagues, visualization evaluation, metrics validation/results, and user bias). More details on methodology, trust levels, and our categorization can be found in our 2020 STAR.
GENERAL OVERVIEW
Time & Venues
Our collection includes 542 techniques sourced from various journals, conferences, and workshops. Figure 1 shows the temporal distribution of these publications, with a stable growth in interest in the topic since 2009 and a remarkable increase evident in 2021, as well as another significant rise in numbers for 2023.
Table 1 outlines the distribution of publication venues. The majority of the techniques were primarily published on visualization venues. Notable exceptions include workshops such as WHI of ICML and special journal issues such as the human-centered explainable AI of ACM TiiS, aiming to engage with the ML community. Despite these efforts, the limited number of publications in other discipline venues may imply that visualization researchers struggle with external outreach. This situation may also suggest that experts from other fields remain largely unaware of the extensive opportunities offered by the visualization field and highlight the importance of visualization literacy.
Note: ‘#’ column shows the number of articles in this survey using a specific dataset; the datasets are also grouped based on this frequency.
Co-authorship Network
We performed co-authorship analysis in Gephi, weighting author nodes by their number of co-authorship edges (Figure 2). The network has 96 weakly connected components, with the ten largest color-coded.
[…]
Datasets
We analyzed publicly accessible datasets from the 542 articles, sorted by frequency and then by recency. Table 2 lists 35 datasets, found in at least four articles.
The most frequently occurring datasets are MNIST, CIFAR-10/100, ImageNet, Wine Quality, Iris Flower, Fashion MNIST, German Credit Risk, Adult Census, 20 Newsgroups, and Breast Cancer. Of these ten datasets, four are about computer vision and are typically used in articles related to DL and neural networks (NNs). Classification is by far the most frequent target variable, followed by regression and clustering. Table 2 contains information about the number of instances and features, as well as the number of classes (if there are any available). The number of articles that used the datasets is visible in the ‘#’ column of Table 2.
Survey Browser
The interactive survey browser’s user interface (UI) consists of (1) a grid showcasing thumbnails of various visualization techniques and (2) an interactive panel that facilitates users to filter based on categories, time, and text (see Figure 3). Clicking a thumbnail reveals details and bibliographic references for that particular technique. At the top of the webpage, links provide access to dialogs containing detailed statistics for the dataset and supplemental materials. We encourage readers to explore the online survey browser and suggest new articles via the "Add entry" dialog.
ML Processing Phase | 542 | |
Pre-processing / Input | 120 | +4% |
In-processing / Model | 156 | +6% |
Post-processing / Output | 363 | -14% |
Treatment Method | 542 | |
Model-agnostic / Black Box | 395 | +1% |
Model-specific / White Box | 165 | -5% |
Dimensionality | 542 | |
2D | 538 | +1% |
3D | 9 | -1% |
Visual Aspects | 542 | |
Computed | 497 | -6% |
Mapped | 273 | -5% |
Visual Granularity | 542 | |
Aggregated Information | 468 | -6% |
Instance-based / Individual | 377 | -3% |
Visual Representation | 542 | |
Bar Charts | 243 | +4% |
Box Plots | 39 | +1% |
Matrix | 153 | +3% |
Glyphs / Icons / Thumbnails | 131 | -8% |
Grid-based Approaches | 59 | +1% |
Heatmaps | 149 | +4% |
Histograms | 174 | +4% |
Icicle Plots | 9 | -1% |
Line Charts | 154 | 0% - |
Node-link Diagrams | 128 | 0% - |
Parallel Coordinates Plots (PCPs) | 73 | -3% |
Pixel-based Approaches | 20 | 0% - |
Radial Layouts | 90 | +6% |
Scatterplot Matrices (SPLOMs) | 26 | -4% |
Scatterplot / Projections | 294 | -4% |
Similarity Layouts | 190 | +21% |
Tables / Lists | 201 | -6% |
Treemaps | 13 | -1% |
Other | 207 | +8% |
Interaction Technique | 525 | |
Select | 491 | +9% |
Explore / Browse | 445 | -3% |
Reconfigure | 220 | +4% |
Encode | 294 | -2% |
Filter / Query | 314 | +1% |
Abstract / Elaborate | 354 | -24% |
Connect | 395 | +9% |
Guide / Sheperd | 169 | +7% |
Verbalize | 31 | +1% |
Note: Each row shows the total count of techniques per category (with heatmap-style icons), and the % difference compared to the 2020 STAR.
IN-DEPTH ANALYSIS
Trust in ML Visualization Revisited
Table 3 shows the most prevalent aspects of existing visualization techniques for increasing trust in ML.
For Data, computer vision ( of 542 articles; compared to the 2020 STAR), health (; ), business (; ), and humanities (; ) are the most prominent domains, while the biology’s popularity . Multi-class classification (; ) is still by far the most frequently found target variable.
For ML methods, non-linear (; ) and then linear DR (; ) techniques are commonly used (the opposite was true in the 2020 STAR), followed by bagging (ensemble learning) () and CNNs () from the DL subcategory (; ). The vast majority of the techniques address supervised learning () and specifically classification problems (; ), and then unsupervised learning (; ) with DR (; ) and clustering (; ).
The ratio of in-processing (; ) to post-processing (; ) techniques has increased from to in the past three years. Most techniques are built as model-agnostic (; ) to target various ML methods and a bigger user audience.
The absolute majority of the visualizations rely solely on 2D representations (; ). For visual aspects and granularity, almost all techniques used have at least a computed (; ) component that is not directly mapped data (; ). Aggregated information (; ) is more common than instance-based/individual instances’ exploration (; ).
Popular visualizations are scatterplots (; ), bar charts (; ), histograms (; ), line charts (), heatmaps (; ), gylphs/icons (; ), and node-link diagrams (). Simpler ones, e.g., tables/lists (; ) and matrices (; ), work better with instance-based exploration.
On the interaction side, selection (; ), exploration (; ), and connection (; ) between different views are the three most common categories found in numerous articles, followed by other interaction techniques, such as abstraction/elaboration (; ), filtering out/searching for (; ) specific instances, and encoding (; ).
Color (; ) is the primary visual channel employed for conveying information in various VA tools and systems. The extensive utilization of opacity (; ) for hiding data points/instances as well as size/area (; ) for representing data variables can be attributed to the widespread use of scatterplots.
As for the evaluation, of the visualization techniques we analyzed have not been evaluated by users (), which is a huge improvement from the three years before. If we consider that the unevaluated visualization techniques could undergo extensive quantitative testing with different (synthetic) benchmarking datasets (e.g., a common approach for DR visualization), this shows a huge shift in how important the visualization community considers evaluation.
For TL1, more works tackle source reliability problems (; ) rather than the transparent collection process challenge (; ). For TL2, researchers focus on the comparison of structures (; ), an increase of prior to the past three years, and then outlier detection (; ) and guidance/recommendations (; ). For TL3, understanding (; ), steering (; ), debugging (; ), and comparing (; ) ML methods are quite popular. Understanding became more common recently, and more techniques are designed for debugging than comparing in the past three years. For TL4, performance (; ), in situ comparison (; ), and what-if hypotheses (; ) are commonly emerging categories associated with a concrete ML model selection. For TL5, metrics validation and results observation is the most frequent category covering of all the articles ().
The visualization techniques have as a main target group, usually practitioners/domain experts (; ), followed by ML experts (; ) with a huge difference. The most underrepresented groups are model developers (; ) and beginners (; ).
In summary, the ML side primarily relies on model performance and metrics validation to increase trust in ML, and the visualization side often opts for scalable multivariate visualizations preferred by experts.
Temporal Trends
We have also analyzed the temporal distribution for each category (normalized per respective year) to find which categories have been more or less prominent in the field in the last three years, as displayed in Figure 4.
These results suggest that business, computer vision, and health data are the most commonly targeted domains. Multi-class data is steadily at the forefront of research while multi-label data is still an underrepresented category with no trend for a potential increase.
For ML methods, we see an increase in other DL methods, such as explaining (vision) transformer NNs, as well as CNNs as the specific NN architecture most prominent compared to all others. Techniques for non-linear DR are more likely to be researched in recent years compared to linear DR methods. Bagging is the most common ensemble learning method. For ML types, we observe a small decline in visualization techniques for clustering and DR, and a subtle increase in semi-supervised learning. A more interesting result here can be observed in the ML processing phase, where in-processing is visualized increasingly most often, then pre-processing, and finally post-processing. Model-agnostic techniques are increasingly more common in the community compared to model-specific. Computed and aggregated information are gradually rising compared to their counterparts.
For the visual representation, the scatterplots and similarity layouts, heatmaps, and radial layouts are more and more commonly used compared to bar charts, matrix representations, glyph/icons, and histograms that remain constant or slightly decrease in popularity. For the interaction techniques, verbalization emerged in 2010 and has not drawn much attention in the visualization community—but recently, there has been some re-appearance compared to the past. Furthermore, opacity, shape, and size are the most usual ways to represent the data after color, while position is almost extinct. Although evaluating visualizations with comparative evaluations is very rare since each technique does not have a suitable direct counterpart, the number of evaluated techniques increases in total compared to the past—expert interviews before, during, or after tool development, but also user studies.
For the TLs, many techniques are not covered much by articles, such as transparent collection processes and source reliability. Notably, equality/data bias started to emerge as a theme in the few articles published in 2023. The visualization community could also explore other intriguing research topics, such as investigating how visualization can assist with the familiarity a user has for a learning method. Comparison of structures, guidance, and outlier detection seem to be in the spotlight based on Figure 4. Additionally, understanding and refinement of learning methods appear to peak in the last year. Knowledgeability about learning methods and details available to diverse user types are slightly better supported than in the past. Moreover, customizable and reconfigurable visualizations that consider the users’ experience level to select a specific ML model remain largely underresearched. A few visualization techniques enable agreement of colleagues and evaluate the use of provenance in VA tools and systems to cover trust in ML issues. User bias is absent from almost all VA systems.
Finally, beginners and model developers are the two least represented target groups.
Category Correlations & Patterns
We have conducted a correlation analysis for our categorization (see Table 3). Individual articles were the instances/observations, and categories were the dimensions/variables. We employed linear correlation analysis to gauge the relationship between category pairs. Given the extensive size of the correlation matrix, we have only provided a small thumbnail representation of it in Table 4(a), with the complete table available in the supplemental material. Table 4(b) shows Pearson’s r coefficient values and highlights distinctive patterns, as well as notable instances of positive (green) and negative (red) correlations among the categories.
In the remainder of this subsection, we focus on the correlations different from our previous survey article published in 2020, but for the sake of completeness, we refer the reader to Table 4 so as to check all interesting correlations found in our dataset’s entirety.
The new findings for negative correlation is that techniques not evaluated by users do not involve visualization interactions, supporting our previous claim that static visualizations are being evaluated via quantitative experiments (if at all). For example, the correlation between linear and non-linear DR with no evaluation is because of the use of synthetic datasets for quantitative experiments. Additionally, beginners rarely use abstract as an interaction technique.
Note: We use red for negative correlations and green for positive ones. Cases highlighted in bold refer to broader groups of categories.
Cases with positive correlation start with an obvious case of a continuous target variable being associated with regression problems. Abstracting further the data leads to more in-depth elaboration with different views that require to be better connected.
Similarity layouts are usually represented with scatterplots compared to other visualizations (e.g., a similarity-based matrix). The comparison of structures is a moderately correlated category with the goal to provide further guidance at the data exploration phase.
Note: The proposed topic titles are in italics. Every topic is encoded by one distinct color. A gray colormap is used for the weight of each term.
Computed data and statistics lead to aggregated information visualization. The model performance is typically tested with metrics validation to choose a concrete ML model. In situ comparison of models occurs most probably for examining model-specific ML algorithms. The comparison of models correlates with knowledgeability and bagging since knowing about how different weak learners work is necessary to use them in an ensemble. Understanding an ML model weakly correlates with techniques for specific models.
Visualizing 2D DR projections with scatterplots is rather usual. Additionally, visualization techniques for pre-processing are correlated with the TLs of raw data and data as a whole. CNNs are the most common ML algorithms for computer vision. Investigating fairness and model bias in ML reassures that models and systems operate equitably without replicating existing biases or introducing new ones, thus making more robust predictions. Model performance in classification problems aims to monitor (and increase) the model accuracy. Guide and refinement visualization loops are important for well-calibrating ML models.
Gaining a deeper understanding and effectively debugging in-processing mechanisms of ML models are central to improving their reliability and usability. Scatterplots are invaluable for outlier detection, offering visual insights into data distributions and aiding in identifying anomalies that might otherwise go unnoticed. Domain experts are uniquely positioned to compare and analyze various structural frameworks, leveraging their specialized knowledge to offer deep insights and critiques. Addressing fairness in the context of social and socioeconomic data tries to ensure that ML technologies do not worsen societal inequalities and disparities. Mapped tables and lists contribute to a more structured and organized representation of data, enhancing accessibility and comprehension for diverse users. For beginners, accumulating hands-on experience is essential in solidifying their understanding and promoting skill development. We also found a weak correlation of shape to design glyphs in visualizations Size is crucial in node-link diagrams, where the dimensions can effectively convey hierarchical relationships and relative importance among interconnected entities.
Furthermore, while the correlation analysis results discussed above focus on pairs of categories, we also conducted frequent pattern mining with our survey data to uncover interactions between multiple categories. Focusing on the patterns with at least 10% support in the data (as the number of possible category combination patterns is otherwise overwhelming), we have detected 12 patterns with 15 categories recurring in our data (the list is provided as part of supplemental materials), while further patterns with 14 or fewer categories could also be considered. The top pattern in the current survey data is supported by 66 articles (12%), and it provides a glimpse into the profile of typical VA approaches in this field: 2D visual representations for computed aspects and aggregated information; support for select, explore/browse, encode, filter/query, abstract/elaborate, and connect tasks; use of color and opacity visual channels; involving the performance aspects, visualization evaluation as well as metrics validation/results, and practitioners/domain experts.
Topic Modeling
Following the same methodology as in our previous articles, we have conducted a topic analysis using latent Dirichlet allocation (LDA) based on the full texts from the PDFs. This approach assigns documents to different clusters with various weights (see supplemental material) and identifies descriptive terms for each cluster. While the topic modeling outcomes are partly influenced by the chosen parameters, they offer insights that complement those obtained from our manual analysis. Thus, the topic analysis confirms and deepens our understanding of the categorized articles.
In our approach, documents get assigned into one of ten clusters according to the topmost weight, with eight descriptive terms per cluster, as shown in Table 5. Based on a thorough review of the top terms and the content of the articles, we manually assigned the topic titles for each topic. Finally, the results are visualized as a DR projection and bar charts (see Figure 5).
Topics
Here, we summarize the ten topics (see Table 5).
T1 This topic comprises 49 articles, focusing on the understanding/explanation of ML models. Multiple visualization tools belonging to this category have been evaluated with user studies involving diverse participants, such as domain and ML experts. Additionally, an unresolved challenge related to this topic is determining the most effective strategies for developing visualization tools that account for ML trustworthiness.
T2 A common theme here is the use of visualization in explaining decision trees (found in 42 articles). Other subtopics are the exploration of behavior regarding the decomposition of instances (or points if DR is used), the user’s role in reasoning with subsets of attributes, and showing how classes are formed in connection to the internal parts of rule-based models.
T3 The 45 articles focus mainly on DL for image data. A recent subtopic here is federated learning, that is, a decentralized training approach allowing distant clients to learn from data while private information is kept hidden. Visualization can help inspect and detect anomalies and errors during training ML algorithms in federated learning. Moreover, this topic is connected to reinforcement learning problems, where more research is needed to study what behaviors are associated with low and high reward levels and to track their evolution throughout the training phase.
T4 This topic has 46 articles allocated, focusing on the use of projections for explaining NN architectures. For example, the recent trend of using transformers for text data—and even extending their use in image applications with the so-called vision transformers—is covered by this topic, with new VA solutions being invented to check how they work.
T5 In DL, significant research efforts have been devoted to understanding the activation of neurons within NNs and visually representing them. Various visualizations, such as 2D saliency and activation maps, have been widely utilized for depicting the activation levels of neurons in diverse DL models, particularly those related to image processing. This topic contains 101 articles that discuss techniques for visualizing the inner workings of NNs during training to detect how they behave in specific instances using projections.
T6 The projection of points with DR is useful for guiding subspace exploration. VA systems are developed for analyzing projection segments to perform error analysis and mine insightful patterns from it. This subtopic and the other about visual analysis of relations between points and dimensions within the various projection spaces were found in 65 articles.
T7 Another area of exploration with 32 articles involves identifying the appropriate distance function. It is essential to ensure that these distances are maintained in the 2D projection generated from the high-dimensional space. This aspect should align with the users’ cognitive expectations of observable clusters.
T8 Incremental data labeling and active learning are a few common themes in this topic of 35 articles. VA systems can suggest which unlabeled instances should be picked first and why to achieve improved predictive performance. Counterfactual explanations are also important for users with systems proposing specific examples that are worthy of investigation.
T9 A prevalent shared aspect among most of the 50 articles in this topic class is their graph NN algorithms and RNNs for text data. Delving into the hidden states of these networks appears to recover lost information that could increase ML trustworthiness under the appropriate guidance of an expert. A hidden subtopic with a few techniques is relevant for the outlier detection catergory that stands out in our categorization with 174 articles in total, as shown in Table 3.
T10 This topic comprises 77 articles that specifically refer to rule-based ML, such as bagging and boosting methods. Rule-based ML models are successful with tabular data and arguably easier for analysts to understand. However, they sometimes reach a level of complexity similar to that of DL, thus it can still be difficult to simplify them.
Compared to our 2020 STAR, Topics 2, 4, 8, 9, & 10 are entirely new due to the recent advancements in graph NNs and (vision) transformers, as well as the need for simpler tree- and rule-based systems for transparent decision-making and ML explanation. The remaining topics are similar but slightly changed compared to the 2020 counterparts. Notably, Topic 1 now refers explicitly to explaining the predictions rather than the ML’s inner workings and evaluating with user studies. Topic 3 is more prominent than before, but hyperparameter tuning is currently disregarded. Explaining individual instances and using DR projections for NN explanation also greatly surged lately (Topic 5).
Topic Embedding
Figure 5(a) presents a 2D projection of the ten-dimensional space of topics. Figure 5(b) reveals that Topics 5 & 10 cover of all articles, followed by Topics 6 & 1, 9 & 3, and the others. Based on Figure 5(c), some interesting top terms are “clusters”, “explanations” (cf. understanding in Table 3), “instances” (instance-based visualization), “trees” (bagging), “participants” (evaluation), “projections” (DR), “layers” (DL), and “image data” (computer vision).
The t-SNE projection in Figure 5(a) shows the purest clusters color-encoded in brown and gray, which are about unique topics (Topics 6 & 8, respectively). Additionally, the misclassification of blue (Topic 1) & cyan (Topic 10) points along with green (Topic 3) & purple (Topic 5) points in the projection occurs because of two conceptual terms shared across these topic pairs, namely, the terms “systems”, “prediction” for 1 & 10 and “NN” and “image” for 3 & 5. Despite Topic 7 being a specific topic, it is associated with the general term: “point”. Thus, a few points in the projection are from Topics 2, 5, 8 & 10 because the individual “points” are related to specific “instances” (or similar terms).
The automatically produced topics further support our categorization. For instance, Topics 1 & 10 represent VA systems focusing on the visualization of prediction explanations with participatory evaluation and prediction performance for rule-based ML, respectively, to facilitate understanding/explanation. Topic 3 is about diagnosing/debugging the reinforcement learning’s training procedure. In contrast, Topics 6 & 7 reflect the comparison of data structures using projections/DR and clustering, respectively. Finally, Topic 2 focuses on the comparison of different tree-based algorithms.
OPEN CHALLENGES
To familiarize readers with each topic, we provide representative references per open challenge (O1–O8).
O1: Improving Popular XAI Methods
Most of the well-known XAI approaches that involve the use of visualization for communicating results were made by experts from other fields. The increasing popularity of these model-agnostic techniques lies in their view of a model as a mathematical function. These functions can be simplified, and interpretation methods can then assess these simpler parts’ behavior. Visualization researchers can contribute to developing VA systems for improving these methods’ interactivity and visual representations, or directly upgrading the techniques’ algorithmic components with their expertise.
O2: New NN Approaches & Self-Supervision
The ML community continually invents new NN architectures or applies existing ones in different settings, while visualization researchers are slow to adopt and work on them. For example, transformer NNs have been used in computer vision settings with promising results compared to convolutional NNs, but still, the visualization community has hardly ever experimented with helping users explore them visually. Similarly, self-supervision is another area of rapid advancement that requires the visualization community’s attention, as well as prompt engineering and model checkers for “hallucinations” of large language models (LLMs).
O3: Multiverse & Confirmatory VA
Most VA solutions, as of recently, were arguably focusing on exploratory visual analysis. There is a trend in utilizing visual representations for confirmatory analysis and hypothesis testing, as well as causal reasoning (e.g., counterfactual reasoning, causal learning, and causal inference), which could be enhanced by visualizations. Another question is how can VA systems guide users through an exhaustive multiverse analysis for choosing the optimal model strategies that will help them understand how a complex model works and decide upon a specific action plan?
O4: Input/Output Uncertainty Quantification
Although the majority of the VA systems aim to help users understand models, quantifying the uncertainty of input and output, as well as checking their robustness and sensitivity to changes, are two major challenges because people want to deploy dynamical, well-calibrated models in real-world practical scenarios. While uncertainty quantification and sensitivity analysis are popular in other fields, visualization research is still limited in checking the inputs and outputs of models, such as with the conformal prediction method that provides statistical guarantees.
O5: Rigorous Evaluation & Benchmarking
Effective collaboration with domain experts from diverse fields often faces substantial communication challenges. These arise from differences in terminology, expectations, and areas of expertise, demanding huge efforts to align objectives and solutions. To mitigate the need for costly user evaluations, the exploration of AI-assisted, simulated evaluation methods is very important. The state-of-the-art benchmarking datasets should also become more rigorous since they may contain non-negligible errors. The development of rational agent benchmarks that estimate the need for and benefits of visualization can also assist XAI methods evaluation. Therefore, such benchmarks potentially minimize the dependency on extensive human participation, thereby reducing associated costs.
O6: Model Deployment & Visual Channels
Scaling VA systems to manage numerous instances, features, and algorithms presents significant challenges, particularly in multi-class classification scenarios. Progressive VA combined with incremental learning is one notable solution, but alternative methods for designing less complex, but more robust systems are essential in real-world settings, where covariate and label shifts are common phenomena. The concurrent use of visual channels, such as color and opacity, for other encoding tasks increases the complexity. Many users find it difficult to navigate these advanced tools, especially at initial exposure. Visualization literacy can partially alleviate this problem by educating people. Additionally, complementary to existing tutorials, storytelling can further illustrate tool functionalities, while the integration of sonification and verbalization can boost user understanding and ease of use of VA systems.
O7: Advancing the Impact & Reproducibility
Many application-specific VA systems are never deployed in the real world, making them susceptible to concept and distribution drifts. This practice highlights the need for further exploration into the out-of-domain/distribution applicability of models, emphasizing the urgency to transform specialized models into more versatile ML solutions. Ensuring the accessibility and usability of VA systems for a broader user base can be achieved through various means, including the integration of visualizations into computational notebooks and offering combined programmatic API and interactive UI solutions that elicit user feedback over long periods of time. Here, the adoption of open-source practices is vital since it not only promotes community involvement and contribution, but also addresses significant challenges related to the replication and reproducibility of ML models used in VA systems.
O8: Underexplored Areas
All underrepresented categories may spark novel research ideas for non-TL categories, implicitly influencing trust in ML. For example, the rareness of visualization tools targeting boosting and stacking ensemble learning methods. Multi-label and regression problems receive less attention than classification. In unsupervised learning, association/pattern mining is less covered by VA tools. Reinforcement learning is also mostly ignored. Finally, the targeted users that require further focus are beginners and people of various experience levels who want to analyze their data in general.
DISCUSSION
A retrospective reflection on the 2020 STAR’s open challenges in relation to Table 3 reveals that the visualization community further researched (a) security vulnerabilities with VA systems for federated learning and adversarial attacks, (b) fairness of the model predictions and counterfactual scenarios, but still without an emphasis on alternative strategic action plans for decision-making (O3), and (c) various forms of biases except for intrinsic visualization and user biases (O5).
Reoccurring challenges in this survey article are O6, O7, and O8, but with further concerns regarding visualization research reproducibility, as well as the deployment of models and VA systems in the wild. Additionally, the demand for approaches fostering visualization literacy and advancing the real-world impact with user-friendly systems that encourage interdisciplinarity collaboration is evident.
Finally, the remaining open challenges (O1, O2, O4) are entirely new and were identified by manually reviewing the content of the categorized articles.
CONCLUSION
In this survey article, we have provided an overview of the updated analyses of the trust in ML visualization techniques dataset maintained by us and provided via an online survey browser. We categorized and conducted various analyses based on the 542 collected, peer-reviewed articles that present a broad range of visualization techniques to improve the trustworthiness of ML models and their outputs. Through our topic, temporal, correlation, and pattern mining analyses, we sought to uncover hidden connections and patterns, as well as identify emerging topics and trends over time for the categorized articles. Our findings reveal the rising enthusiasm for implementing VA tools and systems to enhance trust in ML across diverse data domains, tasks, and interdisciplinary applications. In the future, we plan to keep the online survey browser up to date by continuing to collect and categorize data, especially since LLMs, computer vision, and self-supervision research have recently been thriving with the hope of achieving artificial general intelligence.
ACKNOWLEDGMENTS
This work was partially supported through the ELLIIT environment for strategic research in Sweden.